OpenStudy (anonymous):

Lauren will make annual contributions in the amount of $4,770, on average, to a 401(k) over the next 34 years. Her employer will match 60% of her contributions. She is currently being taxed at 33%, but anticipates being taxed at 15% upon retirement. If her account grows at an average rate of 5.5% annually, what is the value of Lauren’s 401(k) upon retirement? $381,435.69 $395,228.47 $481,057.72 $610,297.11

OpenStudy (anonymous):

Hold on a sec

OpenStudy (anonymous):

ok! :)

OpenStudy (anonymous):

4770(1.60) = $7,632. FVoa = (7632((1.055^34) - 1) / 0.055) = 7632(5.17424 / 0.055) = 7632(94.07712) =$717,006.60 (rounded)

OpenStudy (anonymous):

Did you report that guy or did he just disappear from existence?

OpenStudy (anonymous):

i think he deleted his comment!

OpenStudy (anonymous):

Good choice

OpenStudy (anonymous):

& which answer should i pick because that's not one of my options. :(

OpenStudy (anonymous):

What are your options then?

OpenStudy (anonymous):

$381,435.69 $395,228.47 $481,057.72 $610,297.11

OpenStudy (anonymous):

i really do appreciate your help! :)

OpenStudy (anonymous):

Oh yeah... i'm stupid

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

ughhhh :( idk what to do!

OpenStudy (anonymous):

I'm trying hold on love

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

It's funny that this girl thinks she's gonna retire in 34 years

OpenStudy (anonymous):

lol, i guess

OpenStudy (anonymous):

Wait a second hold on

OpenStudy (anonymous):

alrighty!

OpenStudy (anonymous):

I'm just gonna guess on this one, thank you anyways!

OpenStudy (anonymous):

Wait I think I may be onto something good

OpenStudy (anonymous):

:/ lol okay.

OpenStudy (anonymous):

Ok.... So it turns out I have on sodding idea what the answer is. Your best bet is probably to talk to your teacher because i think you are getting cheated. I checked on yahoo answers and everyone go the same answer, $717,006.60

OpenStudy (anonymous):

aww, well thank you so much for trying!! :') you're awesome.